Describe how the x- and y-coordinates of the given quadratic functions C differ from the x- and y-coordinates of corresponding points of y=x^2

y=-2(x-4)^2
1
Expert's answer
2020-12-14T14:33:38-0500

y = x2

Direction: directed upwards

Top: (0;0)

Focus: (0;1/4).

Symmetry axis: x = 0

Guide: y = -1/4



y =-2(x-4)2

Direction: directed downwards

Top: (4;0)

Focus: (4;−1/8).

Symmetry axis: x = 4

Guide: y = 1/8